Find the value of x.
A 30, 60, 90 degree right triangle is drawn. The 60 degree angle is labeled. The shortest leg adjacent to the 60 degree angle is labeled x. The longest leg opposite the 60 degree angle is labeled 5 times the square root of 3. The hypotenuse is labeled y.

Answer by stanbon(75887)

Find the value of x.
A 30, 60, 90 degree right triangle is drawn. The 60 degree angle is labeled. The shortest leg adjacent to the 60 degree angle is labeled x. The longest leg opposite the 60 degree angle is labeled 5 times the square root of 3. The hypotenuse is labeled y.
-------------
side opposite 30 degree angle = x
---
therefore hypotenuse = y = 2x
---
Pythagoras:
x^2 + (2x)^2 = (5sqrt(3))^2
x^2 + 4x^2 = 75
5x^2 = 75
x^2 = 15
--
x = sqrt(15)
==================
Cheers,
Stan H.
==================